OpenStudy (queenbb):

1. Kim's age is twice that of her sister. When you add Kim's age to her sister's age, you get 36. How old is each sister? (a) Write an equation that represents the situation. Explain any variable used. (b) Solve the equation from Part (a). Show your work. State your solution as a complete sentence. Answer:  

OpenStudy (queenbb):

who ever answers gets a medal

OpenStudy (shadowlegendx):

Let K = Kim's age Let S = Her sisters age \[K = 2S\] \[K + S = 36\]

OpenStudy (shadowlegendx):

Do you understand how I got those equations?

OpenStudy (queenbb):

a little bit are those the answers

OpenStudy (shadowlegendx):

Haha, the above equations and definitions of the variables, answer only part a

OpenStudy (shadowlegendx):

What do you think we should do now, going from here?

OpenStudy (queenbb):

slove it

OpenStudy (shadowlegendx):

How?

OpenStudy (queenbb):

36 -2

OpenStudy (shadowlegendx):

How did you get that?

OpenStudy (queenbb):

becuase you had 2S and then there was 36 so i did 36-2 but now i know its wrong by looking at it

OpenStudy (shadowlegendx):

Ok, we want to solve for K and S, because the question is asking "How old is each sister?"

OpenStudy (queenbb):

umm

OpenStudy (shadowlegendx):

Remember how we defined each variable Let K = Kim's age Let S = Her sisters age Then set up the equations \[K = 2S\] \[K + S = 36\] We have setup a system of equations, comparing the two variables. There are two algebraic methods you can use to solve for this type of problem. Have you heard of elimination or substitution before?

OpenStudy (queenbb):

i think so

OpenStudy (shadowlegendx):

Which one have you heard of, and do you have an idea of how you are supposed to start it?

OpenStudy (queenbb):

no

OpenStudy (shadowlegendx):

Lets go with substitution, as it is the easiest one to learn. Basically, substitution is when you solve for one variable, using the equations, to solve for the other one. First, we want to isolate one variable, and use what it is equal to. We have, \[K = 2S\] and \[K + S = 36\] Since K = 2S, we can substitute 2S, for K, in the equation \[K + S = 36 \rightarrow 2S + S = 36\]

OpenStudy (queenbb):

k

OpenStudy (shadowlegendx):

Add like terms \[3S = 36\] Isolate the variable, S, by dividing the equation by 3 \[S = 12\] Therefore we know that the sisters age is 12. Now we need to solve for Kim's age. Luckily we already have an equation that equals her age \[K = 2S\] We know that, \[S = 12\] Therefore we input... \[K = 2(12)\] Distribute 2, to 12(multiply 2 by 12) \[K = 24\]

OpenStudy (shadowlegendx):

We have, S = 12 and K = 24 Lets check our work using the equations \[K = 2S\] \[24 = 2(12)\] \[24 = 24 \] True \[K + S = 36\] \[24 + 12 = 36\] \[36 = 36\]

OpenStudy (shadowlegendx):

Now we know for sure, that Kim's age is 24, and her sisters age is 12.

OpenStudy (queenbb):

ok so the answer is

OpenStudy (shadowlegendx):

(a) K = Kim's age S = Her sister's age K = 2S K + S = 36 (b) Show the work that I showed you Then use this sentence "Kim's age is 24, and her sisters age is 12."

OpenStudy (shadowlegendx):

Do you understand?

OpenStudy (shadowlegendx):

Follow these steps, I gtg :P